OpenStudy (anonymous):

How do I state the value of k that makes each expression a perfect square trinomial? a) x²+ 10x + k

OpenStudy (anonymous):

This is basically completing the square.

OpenStudy (anonymous):

\[ (x+s)^2 = x^2+\color{red}{2s}x+\color{blue}{s^2} \]Compare that with out equations:\[ x^2+\color{red}{10}x+\color{blue}k \]Notice that \(\color{red}{2s=10}\) and \(\color{blue}{s^2=k}\). Do you think you can solve for \(s\)?

OpenStudy (anonymous):

I'm confused